36x^2+30x-36 is the product of 6x-4, and another. What is the other factor?
1
Expert's answer
2013-09-05T10:53:38-0400
36x^2+30x-36=6(6x^2+5x-6)
6x^2+5x-6=0

D=25+144=169
roots are x1,2=(-5+-13)/12
x1=-18/12=-3/2
x2=8/12=2/3
So, 36x^2+30x-36=6(6x^2+5x-6)=6*6(x+3/2)(x-2/3)=6(2x+3)(3x-2)=3(2x+3)(6x-4)=(6x+9)(6x-4)